# Perfect Edge Domination in $P_{6}$-free Graphs and in Graphs Without Efficient Edge Dominating Sets

### v1

@@ 6 Weighted minimum PED-set and Counting number of PED-sets
    In Substep 4.a of non-black-vertex sub-algorithm (Configuration I), statements (i) and (ii) should each be evaluated independently, taking their corresponding actions if true (each could lead to a different valid $3$-coloring). This adjustment does not affect the algorithm’s complexity.
@@
The hypothesis of Lemma 4.4 is as follows: $G$ is a $P_{6}$-free graph containing a maximal dominating (not necessarily induced) complete bipartite subgraph $K$ (with partitions $R$ and $S$), and $G$ admits a valid $3$-coloring $(B,Y,W)$ such that $|R\cap Y|=|S\cap Y|=1$. This lemma guarantees that:
@@
  • Case 1: $|R|=|S|=1$. Clearly, $V(G)\setminus V(K)\subseteq W$ and $B=\emptyset$. If this unique $3$-coloring is valid, then its corresponding PED-set is a DIM consisting of exactly one edge, namely the unique edge of $K$.
@@
  • Case 3: $|R|=1$ and $|S|>1$. Since $K$ is maximal, we have $R=\{r\}$ and $S=N_{G}(r)$. Clearly, $r\in Y$. By Lemma 4.4, we analyze all possible valid $3$-colorings according to the number of edges in $G[S]$.
    If $G[S]$ contains at least one edge, then the lemma describes a method to find at most two different possible valid $3$-colorings, and this can be done in $O(m+n)$-time.
@@
    If $Y^{\prime}\cup Y^{\prime\prime}$ is not a dissociation set, or $W^{\prime}\cup W^{\prime\prime}$ is not an independent set or $S^{\prime}=S\setminus(W^{\prime}\cup W^{\prime\prime})=\emptyset$, then there is no valid $3$-coloring that satisfies the hypothesis of Lemma 4.4. This check can be performed in $O(n+m)$-time.
    If instead $Y^{\prime}\cup Y^{\prime\prime}$ is a dissociation set, $W^{\prime}\cup W^{\prime\prime}$ is an independent set, and $S^{\prime}\neq\emptyset$, then the task reduces to finding the vertices $s\in S^{\prime}$ such that the corresponding $3$-coloring
@@ 6.2 Detailed explanations of Modification (3)
In this subsection, we describe how to efficiently determine: (i) the number of distinct PED-sets of $H_{\pi}$ that correspond to valid $3$-colorings with $S_{2}\cup X_{2}\subseteq Y$, and (ii) one such PED-set with minimum weight.
Clearly, the value in (i) is obtained by multiplying the numbers of valid $3$-colorings in each connected component $C$ of $H_{\pi}$ where $V(C)\cap(S_{2}\cup X_{2})\subseteq Y$.
We use the following recursive procedure, Two (with $S^{\pi}_{2},X^{\pi}_{2}$, and $Z$ as arguments), to compute this number and find the PED-set for (ii). This procedure returns a number $Qty$, a subset $Z^{\prime}\subseteq Z$, and the weight of the optimal DIMs. Here, $Qty$ is the answer to (i), and when $Qty\geq 1$, $Z^{\prime}$ completes a valid $3$-coloring of $H_{\pi}$ by assigning:
@@
The PED-set corresponding to this $3$-coloring is the answer to (ii). The correctness of Two is due to Lemma 4.8, and its time complexity is O(mn)O(m*n), since there are at most $|Z^{\prime}|+1$ calls to the procedure.
@@ 7 Conclusions and further results
To the best of our knowledge, little is known about counting PED-sets and DIMs in graphs. For the latter, [19] presents a linear-time algorithm for claw-free graphs. Our PED-set counting algorithm (Section 6) can be adapted to count DIMs in cubic time by considering only colorings (d) and (e) in Fig. 2 and Configuration I; whenever a black vertex is found, we skip to the next partial $3$-coloring. This problem is NP-complete for many graph classes (cf. [12, Table 1]).

### v2

@@ 6 Weighted minimum PED-set and Counting number of PED-sets
    In Substep 4.a of non-black-vertex sub-algorithm (Configuration I), statements (i) and (ii) should each be evaluated independently, taking their corresponding actions if true (each could lead to a different valid 33-coloring). This adjustment does not affect the algorithm’s complexity.
@@
The hypothesis of Lemma 4.4 is as follows: $G$ is a $P_{6}$-free graph containing a maximal dominating (not necessarily induced) complete bipartite subgraph $K$ (with partitions $R$ and $S$), and $G$ admits a valid 33-coloring $(B,Y,W)$ such that $|R\cap Y|=|S\cap Y|=1$. This lemma guarantees that:
@@
  • Case 1: $|R|=|S|=1$. Clearly, $V(G)\setminus V(K)\subseteq W$ and $B=\emptyset$. If this unique 33-coloring is valid, then its corresponding PED-set is a DIM consisting of exactly one edge, namely the unique edge of $K$.
@@
  • Case 3: $|R|=1$ and $|S|>1$. Since $K$ is maximal, we have $R=\{r\}$ and $S=N_{G}(r)$. Clearly, $r\in Y$. By Lemma 4.4, we analyze all possible valid 33-colorings according to the number of edges in $G[S]$.
    If $G[S]$ contains at least one edge, then the lemma describes a method to find at most two different possible valid 33-colorings, and this can be done in $O(m+n)$-time.
@@
    If $Y^{\prime}\cup Y^{\prime\prime}$ is not a dissociation set, or $W^{\prime}\cup W^{\prime\prime}$ is not an independent set or $S^{\prime}=S\setminus(W^{\prime}\cup W^{\prime\prime})=\emptyset$, then there is no valid 33-coloring that satisfies the hypothesis of Lemma $4.$ This check can be performed in $O(n+m)$-time.
    If instead $Y^{\prime}\cup Y^{\prime\prime}$ is a dissociation set, $W^{\prime}\cup W^{\prime\prime}$ is an independent set, and $S^{\prime}\neq\emptyset$, then the task reduces to finding the vertices $s\in S^{\prime}$ such that the corresponding 33-coloring
@@ 6.2 Detailed explanations of Modification (3)
In this subsection, we describe how to efficiently determine: (i) the number of distinct PED-sets of $H_{\pi}$ that correspond to valid 33-colorings with $S_{2}\cup X_{2}\subseteq Y$, and (ii) one such PED-set with minimum weight.
Clearly, the value in (i) is obtained by multiplying the numbers of valid 33-colorings in each connected component $C$ of $H_{\pi}$ where $V(C)\cap(S_{2}\cup X_{2})\subseteq Y$.
We use the following recursive procedure, Two (with $S^{\pi}_{2},X^{\pi}_{2}$, and $Z$ as arguments), to compute this number and find the PED-set for (ii). This procedure returns a number QtyQty, a subset $Z^{\prime}\subseteq Z$, and the weight of the optimal DIMs. Here, QtyQty is the answer to (i), and when $Qty\geq 1$, $Z^{\prime}$ completes a valid 33-coloring of $H_{\pi}$ by assigning:
@@
The PED-set corresponding to this 33-coloring is the answer to (ii). The correctness of Two is due to Lemma 4.8, and its time complexity is O(mn)O(m*n), since there are at most $|Z^{\prime}|+1$ calls to the procedure.
@@ 7 Conclusions and further results
To the best of our knowledge, little is known about counting PED-sets and DIMs in graphs. For the latter, [19] presents a linear-time algorithm for claw-free graphs. Our PED-set counting algorithm (Section 6) can be adapted to count DIMs in cubic time by considering only colorings (d) and (e) in Fig. 2 and Configuration I; whenever a black vertex is found, we skip to the next partial 33-coloring. This problem is NP-complete for many graph classes (cf. [12, Table 1]).